-x+3y=10
2x+8y=-6

Answer by TimothyLamb(4379)

-x+3y=10
2x+8y=-6
---
put the system of linear equations into standard form
---
-x+3y=10
2x+8y=-6
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = -7
y = 1
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---